Determine magnitude of the gravitational force that you exert on another person 2.40 m away. Assume you and the other person each have a mass of 63.0 kg.
1
Expert's answer
2012-05-17T09:57:13-0400
F = G*m1*m2/r^2?
m1 = m2 = 63.0 kg
r = 2.40 m
gravitational constant
(G) = 6.67300 × 10^-11 m^3/kg*s^2

So, getting the answer:
F = 6.67300
× 10^-11 * 63.0 * 63.0 / (2.40 * 2.40) = 6.67300 × 10^-11
* 3969.0 / 5.76 =
4598.1140625 × 10^-11 = 4.5981140625 × 10^-8 N =
approx. 46 nN.